OpenStudy (anonymous):

UNCERTAINTIES! In an experiment to measure the density of a steel ball, you measure the diameter of the ball to be 1.76 +/- 0.02 cm and the mass to be 22 +/1 g. Complete the calculations first and then include the effect of experimental uncertainties.

OpenStudy (anonymous):

a) What is the volume of the ball (formula of a sphere: V = 4/3(pi)r^3? I found the volume but I'm confused with the uncertainties...

OpenStudy (anonymous):

V = 4/3(pi)r^3 = 4/3(pi)(0.88)^3 = 2.85 cm^3

OpenStudy (anonymous):

The one thing I'm sure of is finding the maximum and minimum of the measurement so, 1.76's max is 1.78 and its min is 1.74...

OpenStudy (anonymous):

You are on the right track, just calculate the volume values corresponding to those new diameters.

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

calculate the Volume when diameter = 1.78 AND when diameter = 1.74

OpenStudy (anonymous):

ohh okay, after that?

OpenStudy (anonymous):

Now calculate your highest possible value for density and your lowest possible value for density. where density = mass / volume

OpenStudy (anonymous):

yes got that for the second question, but the uncertainties -- I'm not so sure.